# From Competition to Synergy: Unlocking Reinforcement Learning for Subject-Driven Image Generation

### v1

@@ Abstract
Subject-driven image generation models face a fundamental trade-off between identity preservation (fidelity) and prompt adherence (editability). While online reinforcement learning (RL), specifically GPRO, offers a promising solution, we find that a naive application of GRPO leads to competitive degradation, as the simple linear aggregation of rewards with static weights causes conflicting gradient signals and a misalignment with the temporal dynamics of the diffusion process. To overcome these limitations, we propose Customized-GRPO, a novel framework featuring two key innovations: (i) Synergy-Aware Reward Shaping (SARS), a non-linear mechanism that explicitly penalizes conflicted reward signals and amplifies synergistic ones, providing a sharper and more decisive gradient. (ii) Time-Aware Dynamic Weighting (TDW), which aligns the optimization pressure with the model’s temporal dynamics by prioritizing prompt-following in the early, identity preservation in the later. Extensive experiments demonstrate that our method significantly outperforms naive GRPO baselines, successfully mitigating competitive degradation. Our model achieves a superior balance, generating images that both preserve key identity features and accurately adhere to complex textual prompts.
@@ 1 Introduction
In recent years, subject driven image generation, which aims to create customized images that align with both a textual prompt and the specific subjects in reference images, has garnered substantial interest across both academic and industrial communities (Li et al., 2023; Ruiz et al., 2023; Gal et al., 2022; Ye et al., 2023; Zhang et al., 2024; Huang et al., 2025; Wu et al., 2025). Unlike standard text-to-image (T2I) generation (Rombach et al., 2022; Podell et al., 2024; He et al., 2024), which focuses solely on text–image alignment, this task demands a dual capability: maintaining high-fidelity identity preservation from reference images and ensuring accurate prompt adherence in novel contexts. This dual objective introduces a critical generalization challenge: the model must learn a subject representation that is robust enough to preserve its core identity, yet adaptive enough to integrate seamlessly into new contexts as indicated in the prompt. Achieving this balance between identity fidelity and prompt adherence is central to realizing practical personalized generation.
To address this challenge, current research has primarily evolved along two main paradigms: finetuning-based and adapter-based customization. Finetuning-based approaches, such as DreamBooth (Ruiz et al., 2023) and Textual Inversion (Gal et al., 2022), enable subject-specific customization by fine-tuning model parameters using a small set of reference images to learn the specialized representation of single subject. However, these methods are computationally intensive and requires retraining for each new subject, making it impractical for scalable customization. Instead, Adapter-based approaches (Wu et al., 2025; Ye et al., 2023; Tan et al., 2024; Xiao et al., 2025; Kumari et al., 2025) introduce lightweight cross-modal adapters to bridge text and visual representations without requiring per-subject fine-tuning. However, in practice, they often exhibit poor generalization when confronted with complex prompts or novel subjects. To mitigate this, recent methods have explored more generalized strategies, ranging from large-scaled pre-training (Kumari et al., 2025) to sophisticated attention calibration and network designs that attempt to disentangle identity and prompt representations (Kang et al., 2025; He and Yao, 2025; Huang et al., 2025). Nonetheless, essentially, all of the above methods rely on Supervised Fine-Tuning (SFT), which operates as a form of behavior-cloning paradigm that lacks the capacity to generalize beyond the observed examples. This limitation becomes more pronounced when the model encounters novel prompts or unseen subjects, where the training distribution no longer provides reliable guidance.
@@ 2.1 Subject-Driven Generation
To overcome this limitation, a significant body of work has focused on adapter-based methods. A popular approach involves training lightweight adapters to inject visual conditions into the model’s attention layers (Ye et al., 2023; Tan et al., 2024; Wu et al., 2025; Huang et al., 2025; Zhang et al., 2024). More recently, a distinct inference-time paradigm has emerged. These methods (Shin et al., 2025; Kang et al., 2025; He and Yao, 2025) require no additional training and typically operate by reframing the task as a form of guided inpainting or by directly manipulating the model’s attention maps to mitigate concept confusion.
@@ 6.1 Experiment Setup
Our main experiments are conducted on the subject-driven generation model UNO (Wu et al., 2025), a Diffusion Transformer (DiT) based architecture built upon FLUX.1-dev (Andreas Blattmann, 2024). We freeze the pretrained DiT backbone and the image encoder, training only the injected LoRA (Hu et al., 2022) matrices. We provided training details in Appendix A
@@
We conduct a comprehensive evaluation on the widely-used DreamBench (Ruiz et al., 2023). Our method is compared against a suite of state-of-the-art baselines, including both fine-tuning-based methods: Textual Inversion (Gal et al., 2022), DreamBooth (Ruiz et al., 2023) and finetuning-free methods BLIP-Diffusion (Li et al., 2023), ELITE (Wei et al., 2023), SSR-Encoder (Zhang et al., 2024), OminiControl (Tan et al., 2024), OmniGen (Xiao et al., 2025) and FLUX IP-Adatper (team, 2025). Performance is measured using established metrics: CLIP-T (Radford et al., 2021) for prompt adherence, CLIP-I and DINO Score (Hessel et al., 2021; Caron et al., 2021) for identity preservation, and the HPS-v3 (Ma et al., 2025) for text-image alignment and aesthetics.
@@ 6.2 Main Results
Overall, Customized-GRPO improves upon our UNO base model by 6.8%in the DINO score for ID preservation while simultaneously increasing the HPS-v3 score for prompt following and human preference by 5.7%. The concurrent improvement across both ID preservation and prompt following metrics is a key result, providing clear evidence that our method effectively mitigates the competitive degradation observed in naive approaches and successfully learns a more balanced policy.
@@
We also conduct a pairwise human evaluation comparing Customized-GRPO against the base UNO model. As shown in Figure 5, our method is strongly preferred by human annotators, winning in Prompt Following and ID Preservation by decisive margins of 50% vs. 22% and 54% vs. 21%, respectively. In both criteria, our method is preferred more than twice as often as the base model. More detailed information can be found in Appendix A.
@@ 6.4 Analyses
Effect of GRPO versus SFT.
To further validate the efficacy of our online RL approach, we conduct a direct comparison against Supervised Fine-Tuning (SFT). We fine-tune the UNO (base) model using SFT on the same 10k high-quality data curated for our experiments. The results, presented in Table 2, show that SFT yields only marginal or even negative changes compared to the base model. The DINO score sees a negligible increase from 0.760 to 0.762, while the holistic HPS-v3 score actually decreases from 7.87 to 7.43. This outcome is consistent with the understanding of SFT as a form of behavior cloning. Since the base model is already well-trained on a similar data distribution, simply continuing to imitate this static dataset offers no clear path to resolving the complex fidelity-editability trade-off and can even lead to a form of overfitting that degrades overall perceptual quality. In contrast, our Customized-GRPO method demonstrates significant performance gains across both objectives.
@@
We analyze the impact of the synergy coefficient $\alpha$, which controls the strength of the SARS adjustment. As shown in Figure 6 (a), the results reveal a clear optimal balance. When $\alpha$ increase from 0 to 0.5, we abserve a steady improvement on both objectives, demonstrating the effectiveness of SARS in transforming conflicted signals into a productive gradient. However, further increasing $\alpha$ to 0.8 and 1.0 leads to a significant performance degradation. This indicates that an overly large $\alpha$ induces a risk-averse policy that excessively penalizes conflict, thereby failing to achieve a good balance. Based on this analysis , we adopt $\alpha=0.5$ for all subsequent experiments.
@@ A.1 Implementation Details
Our main experiments are conducted on the subject-driven generation model UNO (Wu et al., 2025), a Diffusion Transformer (DiT) based architecture built upon FLUX.1-dev (Andreas Blattmann, 2024). We freeze the pretrained DiT backbone and the image encoder, training only the injected LoRA (Hu et al., 2022) matrices. For training-time sampling, we set T = 25 as the total sampling steps. For GPRO, the model enerates 12 images for each prompt and clips the advantage to the range [1e-5, 1e-5]. We use AdamW as the optimizer with a learning rate of 1e-5 and a weight decay coefficient of 0.0001. All experiments are conducted on 8 GPUs with a per-GPU batch size of 1, for a maximum of 1250 training steps.
@@ C.1 Impact of Reward Model Choice
To further validate the robustness and generalizability of our Customized-GRPO framework, we investigate its performance with a diverse suite of reward models. Beyond our primary DINO-Seg and HPSv3 combination, we also experiment with CLIP-I and CLIP-T as alternative reward signals. Furthermore, inspired by recent work (Huang et al., 2024; Peng et al., 2024; Ku et al., 2023; Hu et al., 2023), we evaluate the use of a powerful Vision-Language Model (VLM) including Gemini-2.5-Pro, Qwen-VL-Max and GPT-4o (Team et al., 2023; Bai et al., 2023; OpenAI, 2024) as an automated evaluator for both ID preservation and prompt following.
@@
C.2 Qualitative Results

### v2

@@ Abstract
Subject-driven image generation models face a fundamental trade-off between identity preservation (fidelity) and prompt adherence (editability). While online reinforcement learning (RL), specifically GRPO, offers a promising solution, we find that a naive application of GRPO leads to competitive degradation, as the simple linear aggregation of rewards with static weights causes conflicting gradient signals and a misalignment with the temporal dynamics of the diffusion process. To overcome these limitations, we propose Customized-GRPO, a novel framework featuring two key innovations: (i) Synergy-Aware Reward Shaping (SARS), a non-linear mechanism that explicitly penalizes conflicted reward signals and amplifies synergistic ones, providing a sharper and more decisive gradient. (ii) Time-Aware Dynamic Weighting (TDW), which aligns the optimization pressure with the model’s temporal dynamics by prioritizing prompt-following in the early, identity preservation in the later. Extensive experiments demonstrate that our method significantly outperforms naive GRPO baselines, successfully mitigating competitive degradation. Our model achieves a superior balance, generating images that both preserve key identity features and accurately adhere to complex textual prompts.111  github.com/Safeoffellow/Customized-GRPO
@@ 1 Introduction
In recent years, subject driven image generation, which aims to create customized images that align with both a textual prompt and the specific subjects in reference images, has garnered substantial interest across both academic and industrial communities (Li et al., 2023; Ruiz et al., 2023; Gal et al., 2022; Ye et al., 2023; Zhang et al., 2024; Huang et al., 2025a; Wu et al., 2025). Unlike standard text-to-image (T2I) generation (Rombach et al., 2022; Podell et al., 2024; He et al., 2024), which focuses solely on text–image alignment, this task demands a dual capability: maintaining high-fidelity identity preservation from reference images and ensuring accurate prompt adherence in novel contexts. This dual objective introduces a critical generalization challenge: the model must learn a subject representation that is robust enough to preserve its core identity, yet adaptive enough to integrate seamlessly into new contexts as indicated in the prompt. Achieving this balance between identity fidelity and prompt adherence is central to realizing practical personalized generation.
To address this challenge, current research has primarily evolved along two main paradigms: finetuning-based and adapter-based customization. Finetuning-based approaches, such as DreamBooth (Ruiz et al., 2023) and Textual Inversion (Gal et al., 2022), enable subject-specific customization by fine-tuning model parameters using a small set of reference images to learn the specialized representation of single subject. However, these methods are computationally intensive and requires retraining for each new subject, making it impractical for scalable customization. Instead, Adapter-based approaches (Wu et al., 2025; Ye et al., 2023; Tan et al., 2024; Xiao et al., 2025; Kumari et al., 2025) introduce lightweight cross-modal adapters to bridge text and visual representations without requiring per-subject fine-tuning. However, in practice, they often exhibit poor generalization when confronted with complex prompts or novel subjects. To mitigate this, recent methods have explored more generalized strategies, ranging from large-scaled pre-training (Kumari et al., 2025) to sophisticated attention calibration and network designs that attempt to disentangle identity and prompt representations (Kang et al., 2025; He and Yao, 2025; Huang et al., 2025a). Nonetheless, essentially, all of the above methods rely on Supervised Fine-Tuning (SFT), which operates as a form of behavior-cloning paradigm that lacks the capacity to generalize beyond the observed examples. This limitation becomes more pronounced when the model encounters novel prompts or unseen subjects, where the training distribution no longer provides reliable guidance.
@@ 2.1 Subject-Driven Generation
To overcome this limitation, a significant body of work has focused on adapter-based methods. A popular approach involves training lightweight adapters to inject visual conditions into the model’s attention layers (Ye et al., 2023; Tan et al., 2024; Wu et al., 2025; Huang et al., 2025a; Zhang et al., 2024). More recently, a distinct inference-time paradigm has emerged. These methods (Shin et al., 2025; Kang et al., 2025; He and Yao, 2025) require no additional training and typically operate by reframing the task as a form of guided inpainting or by directly manipulating the model’s attention maps to mitigate concept confusion.
@@ 6.1 Experiment Setup
Our main experiments are conducted on the subject-driven generation model UNO (Wu et al., 2025), a Diffusion Transformer (DiT) based architecture built upon FLUX.1-dev (Andreas Blattmann, 2024). We freeze the pretrained DiT backbone and the image encoder, training only the injected LoRA (Hu et al., 2022) matrices. We provide training details in Appendix A.
@@
We conduct a comprehensive evaluation on the widely-used DreamBench (Ruiz et al., 2023). Our method is compared against a suite of state-of-the-art baselines, including both fine-tuning-based methods: Textual Inversion (Gal et al., 2022), DreamBooth (Ruiz et al., 2023); finetuning-free methods BLIP-Diffusion (Li et al., 2023), ELITE (Wei et al., 2023), SSR-Encoder (Zhang et al., 2024), OminiControl (Tan et al., 2024), OmniGen (Xiao et al., 2025), FLUX IP-Adatper (team, 2025) and RL methods RPO Miao et al. (2024). Performance is measured using established metrics: CLIP-T (Radford et al., 2021) for prompt adherence, CLIP-I and DINO Score (Hessel et al., 2021; Caron et al., 2021) for identity preservation, and the HPS-v3 (Ma et al., 2025) for text-image alignment and aesthetics.
@@ 6.2 Main Results
Overall, Customized-GRPO improves upon our UNO base model by 6.8% in the DINO score for ID preservation while simultaneously increasing the HPS-v3 score for prompt following and human preference by 5.7%. The concurrent improvement across both ID preservation and prompt following metrics provides clear evidence that our method effectively mitigates the competitive degradation observed in naive approaches and successfully learns a more balanced policy.
We conduct a pairwise human evaluation comparing Customized-GRPO against the base UNO model. As shown in Figure 5, our method is strongly preferred by human annotators, winning in Prompt Following and ID Preservation by decisive margins of 50% vs. 22% and 54% vs. 21%, respectively. In both criteria, our method is preferred more than twice as often as the base model.
@@ 6.3 Ablation Study
[…]
$1.0:0$ 0.832 0.874 0.272 6.39
$0.9:0.1$ 0.821 0.866 0.285 7.27
$0.8:0.2$ 0.814 0.864 0.296 7.68
$\boldsymbol{0.7:0.3}$ 0.812 0.862 0.301 8.32
$0.6:0.4$ 0.795 0.848 0.292 7.88
Table 4: Analysis on the choice of Dynamic Weight. The best performance is marked in bold.
@@ 6.4 Analyses
Comparison with RL and SFT.
To further validate the efficacy of our online RL approach, we conduct a direct comparison against Supervised Fine-Tuning (SFT) and a recent offline preference optimization baseline, RPO Miao et al. (2024). We fine-tune the UNO (base) model using SFT and RPO on the same 10k high-quality data curated for our experiments. The results, presented in Table 2, show that SFT yields only marginal or even negative changes compared to the base model. Since the base model is already well-trained on a similar data distribution, simply continuing to imitate this static dataset offers no clear path to resolving the complex trade-off. Conversely, while offline RPO significantly improves identity fidelity (DINO: 0.852), it incurs severe competitive degradation in text alignment, with the HPS-v3 score dropping drastically to 4.47. This indicates that RPO tends to overfit the reference image, failing to balance conflicting objectives due to its static harmonic mean aggregation and the inherent challenge of constructing balanced offline preference pairs (implementation details in Appendix A). In contrast, our Customized-GRPO method demonstrates significant performance gains across both objectives.
Dynamic Weight Bounds Analysis.
For dynamic weight bounds $w_{max},w_{min}$, we evaluate the impact of different dynamic ranges on model performance while keeping other hyper-parameters fixed. The results are presented in Table 4. Our observations reveal two notable trends: overly aggressive decay ($w_{min}\leq 0.1$) causes semantic collapse in the final denoising stages, where the model tends to overfit the reference image and fails to adhere to textual constraints. Conversely, insufficient dynamic range($w_{max}:w_{mian}=0.6:0.4$) fails to resolve competitive degradation, resulting in drops for both DINO (0.795) and HPS-v3 (7.88) compared to the optimal setting. Based on these results, we adopt $w_{max}:w_{mian}=0.7:0.3$ as the default configuration in this work.
@@
We analyze the impact of the synergy coefficient $\alpha$, which controls the strength of the SARS adjustment. As shown in Figure 6 (a), the results reveal a clear optimal balance. When $\alpha$ increase from 0 to 0.5, we observe a steady improvement on both objectives, demonstrating the effectiveness of SARS in transforming conflicted signals into a productive gradient. However, further increasing $\alpha$ to 0.8 and 1.0 leads to a significant performance degradation. This indicates that an overly large $\alpha$ induces a risk-averse policy that excessively penalizes conflict, thereby failing to achieve a good balance. Based on this analysis , we adopt $\alpha=0.5$ for all subsequent experiments.
@@
Synergy Coef. ($\alpha$) Standard TDW Inverted TDW
$uparrow DINO\uparrow HPS-v3\uparrow$
$\alpha=0.0$ 0.795 7.95 0.605 5.65
$\alpha=0.2$ 0.801 8.04 0.609 5.62
$\boldsymbol{\alpha=0.5}$ 0.812 8.32 0.613 5.58
$\alpha=0.8$ 0.835 7.61 0.616 5.49
$\alpha=1.0$ 0.858 6.73 0.620 5.41
Table 5: Analysis of the interaction between Synergy Coefficient and TDW schedules.
Interaction between SARS and TDW.
To systematically analyse the hyperparameter interactions within our framework, we find that SARS and TDW address competitive degradation from distinct dimensions. We propose that TDW acts as the performance foundation, while SARS acts as the performance amplifier. Consequently, the efficacy of SARS depends on a reasonable TDW schedule. To validate this, we conduct comprehensive experiments on DreamBench, varying $\alpha\in 0.0,0.2,0.5,0.8,1.0$ across both the standard TDW schedule and the Inverted TDW schedule (which reverses the curriculum).
As shown in Table 5, aligning the weighting schedule with diffusion dynamics is the foundation for success. An inverted schedule leads to a performance collapse (DINO $0.6$, HPS-v3 $0.3$) regardless of the $\alpha$ value, showing minimal sensitivity to parameter changes. Under the correct Standard TDW, when $\alpha$ increases from 0 to 0.5, we observe a steady improvement on both objectives. However, further increasing $\alpha$ to 0.8 and 1.0 triggers reward hacking, where the model over-optimizes Identity at the significant expense of Prompt adherence.
@@ 7 Conclusion
Acknowledgment
This work was supported in part by the Ningbo Youth Science and Technology Innovation Leading Talent Program (No. 2025QL059), and the Earth System Big Data Platform of the School of Earth Sciences, Zhejiang University.
@@ A.1 Implementation Details
Our main experiments are conducted on the subject-driven generation model UNO (Wu et al., 2025), a Diffusion Transformer (DiT) based architecture built upon FLUX.1-dev (Andreas Blattmann, 2024). We freeze the pretrained DiT backbone and the image encoder, training only the injected LoRA (Hu et al., 2022) matrices. For training-time sampling, we set T = 25 as the total sampling steps. For GRPO, the model enerates 12 images for each prompt and clips the advantage to the range [1e-5, 1e-5]. We use AdamW as the optimizer with a learning rate of 1e-5 and a weight decay coefficient of 0.0001. All experiments are conducted on 8 NVIDIA A100 GPUs with a batch size of 1, for a maximum of 1250 training steps, resulting in a total of approximately 80 wall-clock hours (i.e., 640 GPU hours).
@@ A.2 Evaluation Details
Regarding the comparison with RPO Miao et al. (2024), given that their original data is not publicly available, we strictly replicate their data construction pipeline to curate a paired preference dataset using our own training data Kumari et al. (2025). Specifically, we employ the frozen UNO base model to generate candidate images conditioned solely on text prompts. Subsequently, we adopt RPO’s core Harmonic Mean reward aggregation mechanism to distinguish between winning and losing samples for preference pair construction. Following RPO protocol, we fine-tune UNO using the composite objective proposed, which combines the Offline DPO Rafailov et al. (2023) loss for preference optimization with a Similarity Loss to ensure identity consistency.
@@ C.1 Impact of Reward Model Choice
To further validate the robustness and generalizability of our Customized-GRPO framework, we investigate its performance with a diverse suite of reward models. Beyond our primary DINO-Seg and HPSv3 combination, we also experiment with CLIP-I and CLIP-T as alternative reward signals. Furthermore, inspired by recent work (Huang et al., 2025b; Peng et al., 2024; Ku et al., 2023; Hu et al., 2023; Bi et al., 2025), we evaluate the use of a powerful Vision-Language Model (VLM) including Gemini-2.5-Pro, Qwen-VL-Max and GPT-4o (Team et al., 2023; Bai et al., 2023; OpenAI, 2024) as an automated evaluator for both ID preservation and prompt following.
@@
Method Training Paradigm Iteration Time (s) Training Time (h) Sampling Time (s) Performance (DINO / HPS-v3)
UNO (Base) SFT 2.32 3 8.1 0.762 / 7.43
Naive-GRPO Online RL 291.28 80 8.1 0.801 / 7.56
Customized-GRPO Online RL + SARS + TDW 311.82 80 8.1 0.812 / 8.32
Table 8: Wall-clock time analysis and computational efficiency comparison evaluated on 8 NVIDIA A100 GPUs.
@@
C.2 Computational Efficiency Analysis
To comprehensively evaluate the computational efficiency of our proposed method, we conduct a wall-clock time analysis using 8 NVIDIA A100 GPUs on the 10k subset of the Syncd dataset. The comparison setup is configured as follows:
  • SFT: Batch size = 8, trained for 2 epochs.
  • GRPO (Naive & Ours): Batch size = 1, Group size = 12, trained for 1 epoch.
The quantitative results regarding training and inference efficiency are summarized in Table 8. Based on these results, we derive two main observations:
Training Efficiency: While RL-based training ($\sim$80h) involves higher computational costs than standard SFT (3h) due to the inherent nature of online exploration and reward evaluation, it yields substantial performance gains across both objectives (DINO score increases from $0.762$ to $0.812$; HPS-v3 increases from $7.43$ to $8.32$). Crucially, compared to the Naive-GRPO baseline, our Customized-GRPO incurs negligible additional overhead. This demonstrates that our core contributions—Synergy-Aware Reward Shaping (SARS) and Time-Aware Dynamic Weighting (TDW)—are computationally lightweight and do not burden the overall RL training process.
Sampling Efficiency: Our method maintains a sampling latency ($8.1s$) identical to the base UNO model and Naive-GRPO. Since our optimization strictly updates model weights without altering the architecture, it introduces zero additional cost during inference.
C.3 Qualitative Results